A dielectric material of relative dielectric constant 6 is now inserted and completely fills the space between the plate capacitor.Calculate the new chargee


1
Expert's answer
2021-07-01T12:50:38-0400

Let the capacitance of the capacitor be C.

After inserting the dielectric material in the capacitor, "(C_1)=kC"

Here, in the question it is given , "k=6"

"(C_1)=6C"

Total energy always be conserve,

So, "\frac{Q_1^2}{2C}=\frac{Q_2^2}{2C_2}"

Now, substituting the values,

"Q_2=\frac{Q_1^2 C_2}{C_1}"

"\Rightarrow Q_2=\frac{Q_1\sqrt{\times 6c}}{\sqrt{c}}"

Total charge always be conserve, so it will not change.